OpenStudy (baileycali):

f(x) = 2x − 6. Solve f−1(x) when x = 2. −5 4 −2 7

OpenStudy (visarika23):

Replace f(x) with y y=2x-6 Now write x in terms of y x=(y+6)/2 Replace x with f-1(x) and y with x f-1(x)=(x+6)/2 put x=2 and u will get 4.